OpenStudy (anonymous):

What is the slope and y-intercept of this line? y = x – 3 A. slope: –3, y-intercept: 1 B. slope: 3, y-intercept: –1 C. slope: 1, y-intercept: –3 D. slope: –1, y-intercept: –3

OpenStudy (anonymous):

y = mx + b where m is slope and b is y intercept.

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

so in that equation what is in the same spot as +b? And in the same spot as m

OpenStudy (anonymous):

still dont understand

OpenStudy (anonymous):

can you please explain it better

OpenStudy (lucaz):

y = mx + b where m is slope and b is y intercept your equations is y = x - 3 or y = (1)x + (-3) m = 1, the slope, b = -3, the y intercept

OpenStudy (anonymous):

thanks

OpenStudy (lucaz):

no problem